OpenStudy (anonymous):

How to find the remainder when 51^203 is divided by 7? (using remainder theorem)

OpenStudy (anonymous):

Try to get the pattern

OpenStudy (anonymous):

51^1 % 7 = 2 51^2 % 7 = ?...proceed u will get a pattern

OpenStudy (anonymous):

@chaitanyabdk \[\text{Mod}\left[51^{203},7\right]=4 \] From Mathematica in case you want to verify your answer.